OpenStudy (anonymous):

Allie has a birthday present for her brother. The length of the box that the present comes in is two times longer than the width of the box. The height of the box is 5 inches and its volume is 160 cubic inches. What is the length and width of the box?

OpenStudy (anonymous):

Start with the equation for the volume of a box, Volume = length*width*height. We know the volume to be 160 in^3, the height to be 5 inches, and we know that the length is two times larger than the width, so we can use that information to rewrite the volume equation as: 160 = 2w*w*5 160/5 = 2w*w (divide both sides by 5) 32 = 2w^2 (w*w is w^2) 32/2 = w^2 (divide both sides by 2) 16 = w^2 (square root both sides) 4 = w We now know the width to be 4 inches. Using the initial parameters, we know the length is twice the width, or 2w. Two times 4 is 8. Your length is 8 inches and your width is 4 inches. To verify: 4*8*5 = 160 :)

OpenStudy (anonymous):

Thanks dude!

OpenStudy (anonymous):

No problem :)